### v1

@@ 1 Introduction
With the rapid pace of scientific advancement, there has been a significant increase in the volume of research publications Bornmann and Mutz (2015); Gao et al. (2024). Nevertheless, it poses considerable challenges for traditional scientific feedback mechanisms Liang et al. (2023). On one hand, it exacerbates the pressure on the peer review process Lee et al. (2013); Björk and Solomon (2013); on the other hand, the disparate quality of these numerous publications can negatively affect the scientific research milieu Kelly et al. (2014); Liu and Shah (2023). Consequently, there is an urgent need for an automated scientific reviewing framework designed to generate constructive reviews with strong evidence supports to help authors improve the caliber of their works Yuan et al. (2022).
@@
However, the task of delivering timely, thorough, and perceptive feedback on research papers is inherently intricate and cognitively demanding Horbach and Halffman (2018). Traditional language models typically struggle to handle such lengthy texts, let alone provide valuable review insights Cohan et al. (2020); Wang et al. (2020). Fortunately, Large Language Models (LLMs) have demonstrated emergent capabilities Wei et al. (2022), which have shown state-of-the-art performance in a wide range of natural language tasks Brown et al. (2020); Touvron et al. (2023); Tan et al. (2024). Further, they have also been strengthened to handle increasingly longer contexts Jiang et al. (2023), facilitating the possibility for automated reviewing Liang et al. (2023); Gao et al. (2024).
Currently, some efforts have been made to explore the capabilities of LLMs for automated paper reviewing. For example, Liu and Shah (2023) and Liang et al. (2023) investigate the potential reliability and credibility of paper reviews generated by LLMs with specially designed prompts. Yet most of these LLMs are tailored for broad and general-purpose applications Wei et al. (2023), so simply prompting LLMs in reviewing papers could output generic comments of less value Liang et al. (2023). Further, certain studies have developed peer review datasets and fine-tuned LLMs to learn the paradigm of paper reviewing Wei et al. (2023); Gao et al. (2024). However, in the supervised fine-tuning (SFT) process, these methods simply utilize a review for a paper that can be biased, partial (see Figure 1) and often formalized in various formats and criteria, which could hinder the potential of LLMs for automated paper reviewing Lin et al. (2023); Gao et al. (2024). Also, they lack a self-correction mechanism when the generated reviews are less appealing.
@@
In the Evaluation module, we fine-tune another Mistral-7B to derive the SEA-E model, which can comprehensively analyze papers and generate high-quality reviews. Given papers that are in PDF format, we parse them into text and  codes, and input their corresponding multiple reviews into SEA-S to generate standardized reviews. The parsed papers, standardized reviews and human-crafted prompts constitute another instruction dataset for SFT, leading to SEA-E.
In the Analysis module, we further introduce a self-correction strategy that promotes SEA to rethink and regenerate more constructive reviews, when the generated reviews are inconsistent with the parsed papers. To measure the inconsistency, we put forward a metric, namely, mismatch score. We also train a regression model SEA-A to estimate scores for the generated reviews. Generally, the larger the scores, the less informative the generated reviews.
@@
Finally, we emphasize that the goal of this paper is to provide informative reviews for authors to polish their papers instead of directly recommending acceptance/rejection on papers.
@@ 4.1 Experimental Details
We compare the following baseline methods, which are divided into two categories: (1) Direct inference with LLMs: We directly use Mistral-7B (M-7B) for inference, guided by $inst_{e}$ to generate reviews in the specified format. (2) SFT methods: From all reviews for each paper in the training set, we randomly select one review as the output for SFT, referred to as Mistral-7B-Random (M-7B-R). In addition, gpt-3.5-turbo is used to standardize all reviews for each paper, which is then used as the output in the SFT stage. We call this method Mistral-7B-GPT-3.5 (M-7B-3.5).
@@ 5 Conclusion
In this paper, we presented SEA, a novel framework for automated paper reviewing based on three modules: Standardization, Evaluation, and Analysis. Specifically, we proposed a new paradigm for constructing a standardized review dataset. Based on this dataset, we can fine-tune the long-context LLM to generate high-quality reviews. Moreover, we proposed a new evaluation metric to measure the consistency between papers and generated reviews. Comprehensive experimental results on eight datasets demonstrate that the SEA framework can generate feedback that aligns with human reviews. We anticipate that the SEA framework will help researchers improve the quality of their work and shed light on the field of automated scientific reviewing.

### v2

@@ 1 Introduction
With the rapid pace of scientific advancement, there has been a significant increase in the volume of research publications Bornmann and Mutz (2015); Gao et al. (2024); Lin et al. (2023a). Nevertheless, it poses considerable challenges for traditional scientific feedback mechanisms Liang et al. (2023). On one hand, it exacerbates the pressure on the peer review process Lee et al. (2013); Björk and Solomon (2013); on the other hand, the disparate quality of these numerous publications can negatively affect the scientific research milieu Kelly et al. (2014); Liu and Shah (2023). Consequently, there is a need for an automated scientific reviewing framework designed to generate constructive reviews with strong evidence supports to help authors improve the caliber of their works Yuan et al. (2022).
@@
However, the task of delivering timely, thorough, and perceptive feedback on research papers is inherently intricate and cognitively demanding Horbach and Halffman (2018). Traditional language models typically struggle to handle such lengthy texts, let alone provide valuable review insights Cohan et al. (2020); Wang et al. (2020). Fortunately, Large Language Models (LLMs) have demonstrated emergent capabilities Wei et al. (2022), which have shown state-of-the-art performance in a wide range of tasks Brown et al. (2020); Touvron et al. (2023); Tan et al. (2024); Sun et al. (2024). Further, they have also been strengthened to handle increasingly longer contexts Jiang et al. (2023), facilitating the possibility for automated reviewing Liang et al. (2023); Gao et al. (2024).
Currently, some efforts have been made to explore the capabilities of LLMs for automated paper reviewing. For example, Liu and Shah (2023) and Liang et al. (2023) investigate the potential reliability and credibility of paper reviews generated by LLMs with specially designed prompts. Yet most of these LLMs are tailored for broad and general-purpose applications Wei et al. (2023), so simply prompting LLMs in reviewing papers could output generic comments of less value Liang et al. (2023). Further, certain studies have developed peer review datasets and fine-tuned LLMs to learn the paradigm of paper reviewing Wei et al. (2023); Gao et al. (2024). However, in the supervised fine-tuning (SFT) process, these methods simply utilize a review for a paper that can be biased, partial (see Figure 1) and often formalized in various formats and criteria, which could hinder the potential of LLMs for automated paper reviewing Lin et al. (2023b); Gao et al. (2024). Also, they lack a self-correction mechanism when the generated reviews are less appealing.
@@
In the Evaluation module, we fine-tune another Mistral-7B to derive the SEA-E model, which can comprehensively analyze papers and generate high-quality reviews. Given papers that are in PDF format, we parse them into text and LaTeX codes, and input their corresponding multiple reviews into SEA-S to generate standardized reviews. The parsed papers, standardized reviews and human-crafted prompts constitute another instruction dataset for SFT, leading to SEA-E.
In the Analysis module, we further introduce a self-correction strategy that promotes SEA to rethink and regenerate more constructive reviews, when the generated reviews are inconsistent with the parsed papers. To measure the inconsistency, we put forward a metric, namely, mismatch score. We also train a regression model SEA-A to estimate scores for the generated reviews.
@@
Finally, it is important to emphasize that the purpose of this paper is not to directly recommend the acceptance/rejection on papers. We anticipate our framework SEA can facilitate timely feedback for researchers, thereby enhancing the quality of their work and enabling them to transition efficiently to subsequent projects.
@@ 4.1 Experimental Details
We compare the following baseline methods, which are divided into two categories: (1) Direct inference with LLMs: We directly use Mistral-7B (M-7B) for inference, guided by $inst_{e}$ to generate reviews in the specified format. (2) SFT methods: From all reviews for each paper in the training set, we randomly select one review as the output for SFT, referred to as Mistral-7B-Random (M-7B-R). Mistral-7B-GPT-3.5 (M-7B-3.5) refers to the method where reviews for each paper are standardized using gpt-3.5-turbo, and these standardized outputs are then applied in the SFT stage. Moreover, REVIEWER2 Gao et al. (2024) is a two-stage review generation framework. Due to time-consuming, we use a smaller test set and compare it with REVIEWER2. The detailed experimental results are provided in the Appendix B.
@@ 4.6 Qualitative Decision Analysis
4.7 Human and GPT evaluation
To further validate the performance of SEA, we supplemented the study with a questionnaire experiment for SEA-E and SEA-S, evaluated by both humans and GPT. The specific content of the questionnaire can be found in Appendix C.
4.7.1 Review Quality (SEA-E)
Referring to previous review quality instruments Van Rooyen et al. (1999), we design 10 questions across different aspects, with each question rated on a scale of 1 to 10, where a higher score indicates better review quality. We then randomly sample 20 papers and evaluate the review quality generated by different models.
Human evaluation.
We invite 20 qualified experts to evaluate reviews generated by various models, with each expert randomly assigned to assess 5 papers. Additionally, we include an extra question in the evaluation process: after anonymizing the model names, experts are asked to select the best review for each paper. The evaluation results from each expert were then aggregated, and the mean scores were calculated.
GPT evaluation.
Given the inherent subjectivity in human evaluations and inspired by the work of  Zheng et al. (2023), we employ gpt-4o-2024-05-13 to score the reviews generated by different models using the same set of 10 questions. Due to the limitation of context length, GPT-4o did not perform the top-1 ranking evaluation.
Table 5: Review quality evaluation by humans and GPT. (‘R2’ refers to REVIEWER2.)
Human Evaluation GPT Evaluation
M-7B-R M-7B-3.5 R2 SEA-E M-7B-R M-7B-3.5 R2 SEA-E
Q1 5.1 6.2 3.8 7.9 5.0 6.5 4.5 7.2
Q2 4.6 5.6 3.9 7.8 4.5 6.3 3.9 7.2
Q3 4.5 5.3 3.6 8.0 5.6 6.6 5.1 7.9
Q4 4.3 5.5 3.5 7.8 2.8 3.8 2.7 5.5
Q5 4.4 5.6 4.0 8.0 4.5 5.4 4.2 7.2
Q6 4.0 5.2 3.5 8.2 3.3 4.0 3.1 6.2
Q7 4.4 5.3 3.8 7.4 3.2 4.4 3.1 5.6
Q8 4.5 5.6 3.8 7.7 6.9 7.7 6.3 8.3
Q9 4.6 5.6 3.6 7.8 8.2 8.5 7.0 9.2
Q10 4.5 5.4 3.6 7.9 4.9 5.9 4.4 7.3
Top-1 0.0 0.1 0.0 0.9 -
From the Table 5, it can be seen that SEA-E performs exceptionally well in both GPT-4o and human evaluations, significantly outperforming other models on ten questions. Additionally, in 90% of the cases, SEA-E was preferred by the experts.
4.7.2 Standardized Content Quality (SEA-S)
We design seven questions to assess the standardized content across different models, using the same scoring criteria for evaluating review quality. Then, we randomly select 10 papers for evaluation.
Human evaluation.
We invite 10 experts to evaluate the standardized content generated by different models, with each expert being randomly assigned 5 papers for assessment.
GPT evaluation.
We adopt the GPT-4o to score the standardised content generated by the different models based on the same questionnaire.
Table 6: Standardized content evaluation by humans and GPT.
Human Evaluation GPT Evaluation
M-7B GPT-3.5 GPT-4 SEA-S M-7B GPT-3.5 GPT-4 SEA-S
Q1 6.0 5.6 8.2 9.2 8.4 9.0 9.0 9.1
Q2 6.7 7.5 7.6 8.3 8.9 8.8 8.7 9.3
Q3 5.3 5.9 7.8 9.1 8.5 8.8 9.1 9.3
Q4 7.3 3.7 7.3 9.0 8.7 7.1 8.0 8.9
Q5 7.5 5.0 7.7 9.1 9.6 9.7 9.4 9.9
Q6 6.7 6.8 6.9 8.1 9.3 9.6 9.4 9.9
Q7 5.1 5.2 8.1 8.8 8.5 9.0 9.0 9.2
Table 6 presents the results from two distinct evaluation methods. The human evaluation emphasizes SEA-S’s strengths, especially in Q4, Q5, and Q6, where it excels in content relevance, logical coherence, and conciseness. While GPT-4o generally assigns higher scores across the board, SEA-S consistently stands out compared to other models.
Based on the results of the questionnaire, the effectiveness of the SEA framework is further confirmed. It demonstrates the ability to generate standardized content, leading to comprehensive and high-quality review feedback.
@@ 5 Conclusion
In this paper, we present SEA, a novel framework for automated paper reviewing. Specifically, we propose a new paradigm for constructing a standardized review dataset. Based on this dataset, we fine-tune an LLM to generate high-quality reviews. Moreover, we propose a new evaluation metric to measure the consistency between papers and generated reviews. Comprehensive experimental results demonstrate that the SEA framework can generate feedback that aligns with human reviews. In summary, we emphasize that the initial motivation for the field of automated paper review stems from the time-consuming and labor-intensive nature of traditional peer reviewing. Automated peer reviewing can provide timely feedback, enhancing research quality and accelerate the progress of scientific development. Therefore, we anticipate that the SEA framework will help researchers improve the quality of their work and shed light on the field of automated scientific reviewing.
@@
Acknowledgments
This work is partially supported by the National Key R&D Program of China (2023YFC3341200) and National Science Foundation of China (NSFC) (62137001).
@@ A.3 SEA-A
Interestingly, we apply the SEA framework to this paper and compare its generated review with the official feedback we receive. We find that the SEA framework aligns with some aspects of the actual review in terms of “Strengths,” “Weaknesses,” and “Questions.” In addition, SEA provides further constructive suggestions for improvement, demonstrating its ability to generate comprehensive and high-quality review comments. The specific generated reviews are shown in Figure 9.
@@ Appendix B Compare with REVIEWER2
Appendix C Human and GPT evaluation
In order to evaluate the review quality generated by SEA-E and the standardized content quality generated by SEA-S, we design a questionnaire consisting of 10 questions and 7 questions, respectively. For the review quality, we add an additional question during the human evaluation phase: selecting the best review from those generated by different models. The specific questionnaire contents are as follows:
SEA-E.
  1. 1.
    Did the reviewer discuss the importance of the research question?
  2. 2.
    Did the reviewer discuss the originality of the paper?
  3. 3.
    Did the reviewer clearly identify the strengths and weaknesses of the method (study design, data collection and data analysis)?
  4. 4.
    Did the reviewer make specific useful comments on the writing, organization, tables and figures of the manuscript?
  5. 5.
    Were the reviewer’s comments constructive?
  6. 6.
    Did the reviewer supply appropriate evidence using examples from the paper to substantiate their comments?
  7. 7.
    Did the reviewer comment on the author’s interpretation of the results?
  8. 8.
    Did the reviewer maintain objectivity and avoid bias in the review?
  9. 9.
    Was the language used by the reviewer professional and appropriate?
  10. 10.
    How would you rate the quality of this review overall?
  11. 11.
    Please select what you consider to be the best review. (The question only in the human evaluation)
SEA-S.
  1. 1.
    Did this review effectively consolidate similar points from the original reviews?
  2. 2.
    Did the review maintain objectivity and avoid bias?
  3. 3.
    Did this review contain a lot of useful information?
  4. 4.
    Did this review contain a lot of repetitive content?
  5. 5.
    Is there any contradiction in this review?
  6. 6.
    Was the language used by the review professional and appropriate?
  7. 7.
    How would you rate the quality of this review overall?
Figure 8: Examples of generated reviews for SEA-E and SEA-EA.
Figure 9: The review generated by applying the SEA framework to this paper.